OpenStudy (anonymous):

another problem i'm having trouble with Show that the function F of x equals the integral from x to 4 times x of 1 over t dt is constant on the interval (0, +∞)

OpenStudy (anonymous):

this is the function

OpenStudy (anonymous):

not really sure where to start on this one

OpenStudy (empty):

Hmmm well maybe you could take its derivative, and since the derivative of a constant is 0, if its derivative is 0 then you know it was a constant.

OpenStudy (anonymous):

okay so then \[\frac{ 1 }{ 4x} - 1/x\]

OpenStudy (anonymous):

no wait that cant be right

OpenStudy (anonymous):

\[((\frac{ 1 }{ 4x })(4)) - (\frac{ 1 }{ x })\]

OpenStudy (anonymous):

\[ \frac{ 4 }{ 4x} - \frac{ 1 }{ x }\]

OpenStudy (anonymous):

the 4 cancel out and we end up with zero , and zero is a constant thus the function is constant .... does that seem right to you?

OpenStudy (empty):

Yeah perfect!

OpenStudy (anonymous):

YAY!!

OpenStudy (empty):

Well not quite because zero is a constant slight note

OpenStudy (anonymous):

thanks man !

OpenStudy (empty):

zero is the derivative of a constant!

OpenStudy (anonymous):

okay that makes sense , yea ... ZERO IS THE DERIVATIVE OF A CONSTANT !!

OpenStudy (empty):

Haha yeah so you got zero for your derivative, booooom done! :P

OpenStudy (anonymous):

boom indeed